### v1

# ReynoldsFlow: Exquisite Flow Estimation via Reynolds Transport Theorem

Abstract
Optical flow is a fundamental technique for motion estimation, widely applied in video stabilization, interpolation, and object tracking. Recent advancements in artificial intelligence (AI) have enabled deep learning models to leverage optical flow as an important feature for motion analysis. However, traditional optical flow methods rely on restrictive assumptions, such as brightness constancy and slow motion constraints, limiting their effectiveness in complex scenes. Deep learning-based approaches require extensive training on large domain-specific datasets, making them computationally demanding. Furthermore, optical flow is typically visualized in the HSV color space, which introduces nonlinear distortions when converted to RGB and is highly sensitive to noise, degrading motion representation accuracy. These limitations inherently constrain the performance of downstream models, potentially hindering object tracking and motion analysis tasks. To address these challenges, we propose Reynolds flow, a novel training-free flow estimation inspired by the Reynolds transport theorem, offering a principled approach to modeling complex motion dynamics. Beyond the conventional HSV-based visualization, denoted ReynoldsFlow, we introduce an alternative representation, ReynoldsFlow+, designed to improve flow visualization. We evaluate ReynoldsFlow and ReynoldsFlow+ across three video-based benchmarks: tiny object detection on UAVDB, infrared object detection on Anti-UAV, and pose estimation on GolfDB. Experimental results demonstrate that networks trained with ReynoldsFlow+ achieve state-of-the-art (SOTA) performance, exhibiting improved robustness and efficiency across all tasks. The code is available at https://github.com/wish44165/ReynoldsFlow.
1 Introduction
Optical flow estimates pixel velocities by analyzing intensity changes between consecutive frames, playing a fundamental role in motion estimation [6, 44, 20, 47], object detection [4, 1, 16, 47, 3, 38], scene reconstruction [47, 29], and video stabilization [41, 45, 22]. Historically, these applications have relied on traditional methods such as Horn-Schunck [8] and Lucas-Kanade [25], which are based on assumptions of local brightness constancy and spatial smoothness. However, these assumptions often break down in the presence of occlusions, fine-scale motion, or complex backgrounds. More recently, deep learning-based models, such as FlowNet [10] and FlowNet2 [17], have been proposed to learn optical flow. Additionally, some of these models directly integrate the learned optical flow into their architecture to warp images as an intermediate step for tasks like video frame interpolation. Although these models achieve impressive results, they typically require extensive training on synthetic datasets such as FlyingChairs and FlyingThings3D [10] and often struggle to generalize to real-world scenarios without further fine-tuning. In addition, the general approach to visualizing optical flow often uses HSV-based visualization, which aligns well with human perception by offering better discrimination power through Hue and Saturation. This makes it ideal for tasks such as color tuning, scene segmentation, and computer graphics, as highlighted in [24]. However, while HSV-based optical flow visualization enhances color saliency, it provides limited value for tasks where motion dynamics and structural variations are more critical than chromatic differences.
To preserve the training-free nature, we propose Reynolds flow, a novel optical flow framework grounded in the Reynolds transport theorem. By reinterpreting optical flow as a transport phenomenon of the light field associated with rigid motion, Reynolds flow models a broader range of transport phenomena, including complex motion dynamics and varying lighting conditions. In addition to the conventional HSV-based visualization, ReynoldsFlow, we introduce an enhanced representation, ReynoldsFlow+, which integrates optical flow magnitude, Reynolds flow magnitude, and frame intensity. This approach effectively distinguishes objects in videos with complex backgrounds and varying ambient lighting conditions, addressing the perceptual ambiguities caused by non-uniform color interpretation in HSV space. As demonstrated in Fig. 1, traditional HSV-based optical flow fails to convey accurate motion features in challenging datasets [9, 18, 27], whereas ReynoldsFlow+ provides a more explicit perceptual representation of motion characteristics. Our contributions are as follows:
1.
We propose Reynolds flow, a training-free optical flow estimation framework derived from the Reynolds transport theorem. We remove the brightness constancy assumption and generalize the traditional optical flow methods. This newly proposed flow is computationally efficient and robust for video-based tasks.
2.
We introduce an enhanced visualization, ReynoldsFlow+, improving motion clarity for downstream tasks.
3.
We perform comprehensive evaluations on real-world datasets, demonstrating SOTA accuracy in UAVDB and Anti-UAV for object detection and in GolfDB for pose estimation compared to the existing methods.
The rest of this paper is organized as follows: Sec. 2 reviews challenges in traditional and deep learning optical flow methods. Sec. 3 introduces Reynolds flow, ReynoldsFlow, and ReynoldsFlow+, detailing the theoretical foundations and innovations. Sec. 4 presents experimental results on three datasets and an in-depth analysis. Finally, Sec. 5 summarizes the paper and outlines future directions.
2 Related Work
2.1 Optical Flow Estimation Methods
Early optical flow methods like Horn-Schunck [15] and Lucas-Kanade [25] laid the foundation for optical flow estimation. Horn-Schunck applied global smoothness but struggled with large displacements and motion boundaries. At the same time, though computationally efficient, Lucas-Kanade was effective only for small, consistent motions and struggled in complex scenes. Later refinements, like Optimal Filter Estimation [35] and hybrid methods combining Lucas-Kanade and Horn-Schunck [8], aimed to balance adaptability and accuracy but still faced challenges with occlusion, noise, and textured backgrounds. Methods like Farneback’s polynomial expansion [12] and Brox’s high-accuracy flow estimation [7] focused on improving motion estimation but were computationally expensive. TV-L1 [46] addressed noise resilience, and SimpleFlow [39] offered a non-iterative solution at the cost of accuracy. RLOF [34] advanced feature tracking but still struggled with high-texture regions. Recent techniques aimed to balance accuracy with computational efficiency. DeepFlow [42] integrated deep matching for complex motions, while PCAFlow [43] reduced computational load. The Dense Inverse Search (DIS) algorithm [21] optimized dense flow estimation for time-sensitive applications. Despite these advancements, many methods still struggle with multiscale motion tracking and real-time processing, underscoring the need for improved robustness and efficiency in optical flow estimation.
The advent of deep learning has transformed optical flow estimation. FlowNet [10], the first CNN-based method, greatly improved motion recognition but struggled with fine-grained movements. FlowNet2 [17] enhanced precision with a cascaded architecture, while EpicFlow [32] integrated deep feature matching with variational refinement to improve boundary precision. PWC-Net [36] introduced a pyramid, warping, and cost-volume architecture that improved multiscale motion capture. Similarly, SpyNet [31] enhanced computational efficiency for real-time applications, and RAFT [40] set new benchmarks iteratively refining the flow field using contextual information. However, deep learning-based methods are limited by the need for retraining, posing challenges for deployment in resource-constrained environments like embedded systems.
2.2 Optical Flow in Video Processing
Optical flow estimation has been extensively studied in the context of standard benchmark datasets such as Sintel [2], KITTI [14], FlyingChairs, and FlyingThings3D [10]. While these datasets serve as valuable evaluation tools for assessing the accuracy of flow estimation methods, they primarily consist of synthetic or short video sequences designed for benchmarking rather than real-world applications. Much of the research focuses on optimizing pixel-wise flow accuracy, often treating optical flow as the end goal, as demonstrated by [42, 43, 10, 17, 32], rather than as a means of addressing practical computer vision challenges. Beyond benchmarking, optical flow has proven effective in practical applications such as video stabilization, where flow fields correct camera motion to produce smoother sequences [23], and frame interpolation, where flow-guided warping enhances temporal resolution for higher frame rates [13]. It also plays a crucial role in motion tracking, supporting object detection and activity recognition [37, 30]. However, optical flow images are typically treated as intermediate data within these frameworks, and the reason likely stems from traditional HSV-based visualizations, which fail to convey detailed texture information as presented in [5], even for large objects, thus restricting their practical application. A task similar to our UAV detection scenario is explored in [38], where a convolutional layer emulating the Lucas-Kanade optical flow is integrated into the YOLO architecture to enhance UAV detection. Similarly, [26] employs the Farneback optical flow method to track a player’s body posture and golf club trajectory, classifying them using multiple feature classification techniques. However, both approaches are ineffective due to the ambiguous visualization of optical flow images. To our knowledge, no existing method has demonstrated significant benefits in detecting moving objects and estimating poses by directly leveraging optical flow images in neural networks. This work investigates optical flow-based features, specifically ReynoldsFlow and ReynoldsFlow+, to analyze the impact on neural network performance in real-world scenarios.
3 Methodology
This section begins with the preliminaries, followed by the formulation of Reynolds flow, using the Reynolds transport theorem to provide a novel perspective on optical flow estimation. We then introduce an RGB-based visualization that enhances the clarity of motion features. Finally, we conduct a runtime analysis, comparing Reynolds flow with existing approaches to demonstrate its efficiency.
3.1 Preliminaries
3.1.1 Helmholtz decomposition
Let be a continuously differentiable vector field defined on a domain . According to the Helmholtz decomposition theorem, can be uniquely expressed as the sum of an irrotational (curl-free) component and a solenoidal (divergence-free) component:
where satifies and satisfies .
3.1.2 Reynolds transport theorem
Consider a smooth function where is defined on a time-dependent domain . A grayscale video can be represented by where is the intensity value at the pixel location in the view of field of the camera and denotes the th-frame image in the video footage. The Reynolds transport theorem shows that:
where is the velocity vector of at time . For video footage, if the total intensity value remains constant and the velocity field is divergent-free for all time, then the Eq. 2 becomes
Assuming the integrand in the Eq. 3 equals zero, we obtain the brightness constancy constraint in the derivation of the traditional optical flows.
3.1.3 Area Jacobian in the domain transformation
Consider a region evolving over time under the influence of a vector field . Let at time , the corresponding vector field and time increment be uniform. Using the explicit Euler approximation, the domain transformation from to can be written as following:
From Eq. 4, the differential wedge product at can be approximated by the following:
Using the fact that and , we have
For sufficiently small , this simplifies to:
As a result, is the approximal Jacobian between area differentials in and in . In other word, we have
3.2 Derivation of Reynolds Flow
Let be the traditional optical flow that satisfies the constancy brightness assumption:
where is piecewise constant on on each window patch of the image. We have
which forms the basis for classical optical flow methods, such as the Lucas-Kanade or Horn-Schunck approach. To extend optical flow estimation beyond traditional formulations, we consider a broader perspective based on the Reynolds transport theorem, which removes the brightness constancy assumption and divergence-free restriction of the vector field . By Eq. 1, the Reynolds transport theorem can be reformulated as:
By applying the Euler method on Eq. 8. the left-hand side (LHS) of the Eq. 8 satisfies
Next, applying the Taylor approximation
and the Helmholtz decomposition of the vector field ,
And the right hand (RHS) of Eq. 8 is simply as following
By equating the LHS and the RHS above and setting , we obtain
The integration by part of the above equation leads to
As usual, we can compute the vector field from Eq. 9 by assuming it remains constant within each local window patch . Specifically, we compute on a window patch, denoted as . To approximate the boundary integral term in Eq. 9, we apply Simpson’s rule:
here
and
The domain integral term in Eq. 9 is approximated as:
here
and
From Eq. 9, Eq. 10 and Eq. 11, we obtain the irrotational flow field:
Recall that the irrotational vector field complements the optical flow field derived from the Reynolds transport theorem. We refer to as the Reynolds flow. To ensure smoothness, we then define:
where represents a Gaussian smoothing operator.
In the derivation of Reynolds flow, the optical flow component naturally cancels out from both sides of Eq. 9, allowing it to capture residual flow caused by lighting variations (e.g., shadows, infrared dimming) and non-rigid motion within the camera’s field of view.
3.3 Flow Representation in Color Spaces
Optical flow visualization typically encodes motion magnitude and direction in HSV color space. The flow must be transformed back to RGB space to visualize it. This transformation is highly nonlinear and may lead to perceptual inconsistencies. Minor variations in hue and saturation may cause significant color shifts, making motion interpretation ambiguous. Moreover, accurate flow estimation remains challenging in scenes with mono textures, complex lighting, and dynamic motion. Tasks like moving object detection and motion analysis can become problematic due to non-uniform color variation when neural networks rely solely on HSV-based optical flow features. The robustness of neural networks can, therefore, deteriorate.
Here, we propose an alternative flow representation for Reynolds flow, denoted by ReynoldsFlow+, in which the traditional optical flow , the Reynolds flow , and the current frame intensity are included. The ReynoldsFlow, , represents the combined flow field encoded in the HSV space as in the traditional optical flow approach. On the other hand, the ReynoldsFlow+, , denotes the introduced flow representation that stacks the magnitudes of the optical flow, Reynolds flow, and frame intensity together, i.e., . Notice that the optical and Reynolds flow velocities’ directional information is unneeded in the ReynoldsFlow+. Moreover, in the ReynoldsFlow+, the red and green channels capture the motion speed and the illuminant variations due to ambient lighting and motion, respectively; meanwhile, the spatial details of the current frame are preserved in the blue channel to enhance the contrast of the flow against the background. As shown in Fig. 1, ReynoldsFlow+ significantly improves the clarity of objects across all datasets.
3.4 Optical Flow Runtime Comparison
We present the runtime comparison of various optical flow algorithms on a system with an Intel Core i7-12650H CPU, NVIDIA RTX 4050 GPU, and 16 GB RAM. Specifically, we measure the average processing time per image on UAVDB using either CPU- or GPU-based implementations using OpenCV’s publicly available packages. The GPU-accelerated methods, with the ”cuda_” prefix in the table, include Lucas-Kanade, Farneback, Brox, and TV-L1. Though a direct comparison across all methods is somewhat unfair, as they are not executed on the same processor type, we still select each method’s fastest available OpenCV implementation to provide a realistic evaluation. For instance, we use the GPU-based Farneback method instead of its CPU counterpart due to its significantly higher efficiency. Similarly, other methods follow such selection criteria. The results in Tab. 1 demonstrate that ReynoldsFlow and ReynoldsFlow+ achieve the lowest computation times among CPU-based methods. In particular, ReynoldsFlow+ is slightly faster than ReynoldsFlow, eliminating the need for angle computation and HSV to RGB conversion. Consequently, ReynoldsFlow+ achieves the most efficient runtime and is almost comparable to the existing GPU-based approaches. Furthermore, our method is inherently adaptable to GPU acceleration, making ReynoldsFlow+ an efficient and practical choice for real-time applications.
Algorithms OpenCV Packages Runtime Horn-Schunck [15] – 1.951 Lucas-Kanade [25] cuda_DensePyrLKOpticalFlow 0.013 Farneback [12] cuda_FarnebackOpticalFlow 0.031 Brox [7] cuda_BroxOpticalFlow 0.093 TV-L1 [46] cuda_OpticalFlowDual_TVL1 3.165 SimpleFlow [39] createOptFlow_SimpleFlow 3.302 RLOF [34] createOptFlow_DenseRLOF 0.306 DeepFlow [42] createOptFlow_DeepFlow 2.521 PCAFlow [43] createOptFlow_PCAFlow 0.403 DIS [21] DISOpticalFlow_create 0.046 ReynoldsFlow – 0.042 ReynoldsFlow+ – 0.041
4 Experimental Results
To evaluate the proposed ReynoldsFlow and ReynoldsFlow+, we conducted experiments on two tasks: 1) object detection on UAVDB [9] and Anti-UAV [18] and 2) pose estimation on GolfDB [27]. Compared against 12 approaches: 1) RGB videos, 2) grayscale videos and optical flow methods, 3) Horn-Schunck, 4) dense Lucas-Kanade, 5) Farneback, 6) Brox, 7) TV-L1, 8) SimpleFlow, 9) RLOF, 10) DeepFlow, 11) PCAFlow, and 12) DIS. These methods were selected for their training-free nature, which aligns with our approach. All evaluations in Sec. 4 were performed on a high-performance computing system [28] with an NVIDIA H100 GPU with 80 GB memory.
4.1 Object Detection on UAVDB and Anti-UAV
We evaluate the proposed method in object detection tasks on two distinct datasets: UAVDB and Anti-UAV. UAVDB consists of high-resolution RGB frames (19201080 to 38402160) captured by a fixed camera, while Anti-UAV contains low-resolution infrared frames (512512 to 640512) captured by a moving camera. Both datasets feature single-class annotations and include multiple scales of UAVs, reflecting real-world scenarios where object sizes vary based on their distance from the camera. The UAVDB dataset includes 10,763 training images, 2,720 validation images, and 4,578 test images. Moreover, in the Anti-UAV dataset, we selected 4,800 training images, 1,600 validation images, and 1,600 test images from the 223 videos.
In the implementation details, we use YOLOv11n [19] as the detector, and all models were trained with eight workers, an input image size of 640640, a batch size of 64, and over 100 epochs. Mosaic augmentation was applied throughout the training process, except during the last ten epochs, to stabilize performance. To leverage prior knowledge, we employed transfer learning by initializing the models with official YOLOv11n pre-trained weights and fine-tuning them on the UAVDB and Anti-UAV datasets. For evaluation, we report and metrics on the validation and test datasets. As shown in Tab. 2, YOLOv11n with ReynoldsFlow+ input achieves the highest performance in both datasets, demonstrating its effectiveness. Hereafter, we refer to this configuration using ReynoldsFlow+ input as YOLOv11nRF+.
Methods UAVDB Anti-UAV YOLOv11n [19] RGB 0.849 0.549 0.811 0.518 – Grayscale / Infrared 0.770 0.338 0.660 0.281 0.785 0.424 0.781 0.418 Horn-Schunck [15] 0.062 0.021 0.104 0.021 0.212 0.084 0.217 0.080 Lucas-Kanade [25] 0.253 0.152 0.500 0.200 0.276 0.110 0.280 0.114 Farneback [12] 0.465 0.264 0.258 0.145 0.476 0.235 0.500 0.246 Brox [7] 0.454 0.218 0.244 0.110 0.396 0.172 0.388 0.168 TV-L1 [46] 0.719 0.364 0.779 0.409 0.567 0.260 0.600 0.278 SimpleFlow [39] 0.566 0.338 0.444 0.250 0.118 0.047 0.116 0.049 RLOF [34] 0.172 0.079 0.097 0.039 0.477 0.212 0.472 0.222 DeepFlow [42] 0.118 0.050 0.154 0.058 0.372 0.153 0.344 0.143 PCAFlow [43] 0.399 0.235 0.547 0.332 0.580 0.278 0.580 0.286 DIS [21] 0.345 0.149 0.151 0.057 0.335 0.145 0.347 0.144 ReynoldsFlow 0.500 0.325 0.500 0.288 0.643 0.313 0.646 0.320 ReynoldsFlow+ 0.910 0.571 0.895 0.547 0.812 0.465 0.792 0.446
4.2 Pose Estimation on GolfDB
We then evaluate the proposed method on the pose estimation task on GolfDB, which consists of 1,400 videos at a resolution of 160160 pixels, with each video featuring subjects that occupy nearly the entire frame. The goal is to accurately identify specific poses within the golf swing sequence, which is divided into eight distinct events: Address (A), Toe-up (TU), Mid-backswing (MB), Top (T), Mid-downswing (MD), Impact (I), Mid-follow-through (MFT), and Finish (F). The dataset includes face-on and down-the-line views, offering diverse perspectives to capture the nuanced transitions between each event.
In the detailed implementation, we use SwingNet [27], a model designed for pose estimation in golf swing videos. We maintain the original training configuration, including a sequence length of 64, 10 frozen layers, a batch size of 22, 2,000 iterations, and six workers. We use the pre-trained MobileNetV2 model provided by [33] and then fine-tune it on the GolfDB dataset. Performance is evaluated using the Percentage of Correct Events (PCE) metric, with four-fold cross-validation to ensure reliability. For real-time videos sampled at 30 fps, a tolerance of is used. For slow-motion videos, the tolerance is calculated as , where is the number of frames from Address to Impact, is the sampling frequency, and denotes rounding to the nearest integer. As shown in Tab. 3, SwingNet with the ReynoldsFlow+ input achieves higher PCE than the other configurations. This configuration is referred to as SwingNetRF+.
Methods Average Loss PCE SwingNet [27] RGB 0.2313 0.705 Grayscale 0.2476 0.698 Horn-Schunck [15] 0.2557 0.707 Lucas-Kanade [25] 0.2268 0.692 Farneback [12] 0.2545 0.717 Brox [7] 0.2422 0.708 TV-L1 [46] 0.1769 0.810 SimpleFlow [39] 0.2775 0.665 RLOF [34] 0.2562 0.668 DeepFlow [42] 0.2142 0.716 PCAFlow [43] 0.1885 0.765 DIS [21] 0.1995 0.772 ReynoldsFlow 0.1763 0.804 ReynoldsFlow+ 0.1857 0.812
4.3 Effectiveness of ReynoldsFlow+
To further assess the effectiveness of ReynoldsFlow+, we utilize HiResCAM [11] to visualize the activation maps learned by YOLOv11nRF+. The resulting heatmaps highlight regions of interest, revealing the model’s focus on key structures or objects. As shown in Fig. 2, the activation regions in YOLOv11nRF+ strongly align with the motion cues in ReynoldsFlow+ images for both UAVDB and Anti-UAV datasets. In contrast, the original YOLOv11n exhibits weak or inconsistent activation. These results suggest that ReynoldsFlow+ enhances the model’s ability to capture motion dynamics, reinforcing its significance for UAV detection. Furthermore, the findings highlight the critical role of input modalities in shaping the model’s decision-making process and provide deeper interpretability of motion analysis through HiResCAM. Heatmaps are unavailable for the GolfDB test since SwingNet employs a bidirectional LSTM for event probability prediction instead of relying on convolutional features. To evaluate the impact of ReynoldsFlow+, we analyze probability curves over time for each pose event. Fig. 3 (a) presents the keyframes of the eight events alongside their corresponding ReynoldsFlow+ images. In Fig. 3 (b), the solid and dashed lines denote the probability curves obtained using the original SwingNet and SwingNetRF+, respectively. Ground-truth event times are indicated by dashed vertical gray lines with corresponding event labels and the exact frame number. The results demonstrate that SwingNetRF+ consistently produces higher probability peaks across all events than the original SwingNet, significantly reducing prediction errors and uncertainty. Moreover, SwingNetRF+ achieves precise event frame predictions across all stages. These findings confirm that ReynoldsFlow+, by capturing motion dynamics, substantially enhances event detection accuracy in SwingNet.
Our experiments demonstrate that, except for ReynoldsFlow+, using optical flow images as input failed to outperform YOLOv11n with original video inputs on the UAVDB and Anti-UAV datasets. This suggests that directional motion features in traditional optical flow may be less crucial for UAV detection, as UAVs can move in arbitrary directions. In contrast, while some conventional optical flow methods using HSV-based visualization improved performance on GolfDB, where motion direction seems more intuitive and significant in pose estimation, ReynoldsFlow+ consistently outperformed these without relying on directional information. This is further supported by its more informative heatmaps and probability curves. Overall, ReynoldsFlow+ proves highly effective in detecting small objects and achieves the highest pose estimation accuracy.
5 Conclusion
This work proposes Reynolds flow, a training-free optical flow estimation framework grounded in the Reynolds transport theorem that generalizes traditional methods. Building on this framework, we introduce ReynoldsFlow and ReynoldsFlow+. We train YOLOv11n using these flow videos. Our experiments demonstrate that ReynoldsFlow+ provides more informative motion features even without the direction information of the flows. We evaluated these methods on diverse tasks, including object detection on UAVDB and Anti-UAV and pose estimation on GolfDB. ReynoldsFlow+ consistently improves accuracy on varying scales and complexities, serving as a practical preprocessing module that mitigates downstream learning difficulty without additional computational overhead. In addition, its high computational efficiency makes ReynoldsFlow+ ideal for real-time deployment on resource-constrained edge devices. Future work will explore the integration of ReynoldsFlow+ with learning-based systems on FPGAs to enhance scalability and versatility further.